OpenStudy (anonymous):

Let f(x)={ x+1, x<1 x^2-3x+4, 1less or equal to x and less/equal3 5-x, x>3 find all values of x where f(x) is discontinous

OpenStudy (sirm3d):

\[\Large f(x)=\begin{cases} x+1 & ,x<1\\x^2-3x+4&,1\leq x\leq 3\\5-x &,x>3\end{cases}\] evaluate \[\Large \lim_{x\rightarrow a} f(x)\] where \(a\) is a breakpoint in the intervals, that is, \(a=1,a=3\)